OpenStudy (anonymous):

please help evaluate -6+-4 / 4(8--4^2)

OpenStudy (anonymous):

+- = - -- = + So, -6+-4 / 4(8--4^2) = -6 - 4 / 4(8+4^2) = ...?

OpenStudy (anonymous):

-10/ 48 this is what i got

OpenStudy (anonymous):

Not really, if I haven't read the question wrong. 8+4^2 =...?